Apply weierstrass' inequality to prove that
[nΣi=1 (1/√i)] ≤ 1/√n! [nΠi=2 (√i-1)] +2[nΣi=2 (1/√i)]
1
Expert's answer
2020-12-14T18:37:45-0500

"\u03a3^n_{i=1} (1\/\u221ai)\u2264 1\/\u221an! [\u03a0^n_{i=2} (\u221ai-1)] +2[\u03a3^n_{i=2} (1\/\u221ai)]"

For i=2:

"1+1/\sqrt{2}=(1/\sqrt2)(\sqrt2-1)+2/\sqrt{2}"


Weierstrass' inequality:

"\Pi^n_{i=1}(1-a_i)\geq1-\Sigma^n_{i=1}a_i"

"a_i\in[0..1]"


Let "a_i=1/\sqrt{i}"

Then:

"\u03a3^n_{i=1}a_i\u2264 1\/\u221an! [\u03a0^n_{i=2} (1\/a_i-1)] +2\u03a3^n_{i=2}a_i"

"1≤ 1/√n! [Π^n_{i=2} (\frac{1-a_i}{a_i})] +Σ^n_{i=2}a_i"

"1-Σ^n_{i=2}a_i≤ \frac{1}{\sqrt{n!}} Π^n_{i=2} (\frac{1-a_i}{a_i})"

But, since

"\frac{1}{\sqrt{n!}\Pi^n_{i=2}a_i}=\frac{\Pi^n_{i=2}\sqrt{i}}{\sqrt{n!}}=1"


we get Weierstrass' inequality.